### v1

# KAAE: Numerical Reasoning for Knowledge Graphs via Knowledge-aware Attributes Learning

Abstract.
Numerical reasoning is pivotal in various artificial intelligence applications, such as natural language processing and recommender systems, where it involves using entities, relations, and attribute values (e.g., weight, length) to infer new factual relations (e.g., the Nile is longer than the Amazon). However, existing approaches encounter two critical challenges in modeling: (1) semantic relevance—the challenge of insufficiently capturing the necessary contextual interactions among entities, relations, and numerical attributes, often resulting in suboptimal inference; and (2) semantic ambiguity—the difficulty in accurately distinguishing ordinal relationships during numerical reasoning, which compromises the generation of high-quality samples and limits the effectiveness of contrastive learning. To address these challenges, we propose the novel Knowledge-Aware Attributes Embedding model (KAAE) for knowledge graph embeddings in numerical reasoning. Specifically, to overcome the challenge of semantic relevance, we introduce a Mixture-of-Experts-Knowledge-Aware (MoEKA) Encoder, designed to integrate the semantics of entities, relations, and numerical attributes into a joint semantic space. To tackle semantic ambiguity, we implement a new ordinal knowledge contrastive learning (OKCL) strategy that generates high-quality ordinal samples from the original data with the aid of ordinal relations, capturing fine-grained semantic nuances essential for accurate numerical reasoning. Experiments on three public benchmark datasets demonstrate the superior performance of KAAE across various attribute value distributions.
Numerical Reasoning, Knowledge Aware, Knowledge Graph
ccs: Computing methodologies Knowledge representation and reasoning
Figure 1. An example of KG-based numerical reasoning. Dashed lines indicate facts that are not relevant to the query, solid lines and underlined attributes indicate relevant facts.
1. Introduction
Knowledge Graphs (KGs) serve as structured representations of real-world facts and knowledge, playing a pivotal role in powering a wide array of downstream applications, such as recommender systems (RS) (Cao et al., 2022a; Li et al., 2020; Huang et al., 2019), natural language processing (NLP) (Guo et al., 2023; Lu et al., 2023; Sun et al., 2023) and multimodal tasks(Wang et al., 2023; Li et al., 2023; Ni et al., 2023). Among them, numerical reasoning within KGs emerges as a crucial capability, aimed at identifying and inferring missing numerical relations or hierarchical orders among entities.
Existing methods typically integrate additional attributes of entities into traditional knowledge graph embedding (KGE) models, leveraging entities with rich attribute information for numerical reasoning tasks. Others embed numerical entities directly into a vector space or use Graph Neural Networks (GNNs) to enhance KGE performance. However, these approaches often face critical limitations:
  1. C1
    Semantic relevance: Most current models fail to capture the intricate interactions among entities, relations, and numerical attributes sufficiently. This deficiency results in suboptimal inference outcomes, as these models do not account for how the significance of numeric attributes changes depending on the relational context. For example, in addressing the query (LeBron James, Taller_Than, ?), the height attribute is crucial, whereas for (LeBron James, Younger_Than, ?), age becomes relevant. This selective relevance is often overlooked, leading to the underutilization of semantic details inherent in the numeric attributes, as depicted in Figure 1.
  2. C2
    Semantic ambiguity: Existing models struggle with accurately distinguishing ordinal relationships between numerical values, which are vital for high-quality inference. Numerical values close in magnitude (e.g., 183cm and 184cm) often carry similar semantic meanings but are prone to erroneous ordinal interpretations under current systems. Such ambiguities can significantly impair the model’s ability to perform effective numerical reasoning.
Given these considerations, we pose the following question: Can we develop a novel KGE model that can conduct knowledge-aware attributes learning and exploit latent numerical information within a unified framework?
To address these challenges, we introduce a new model termed KAAE (Knowledge-Aware Attributes Embedding), crafted for numerical reasoning via knowledge graphs. Specifically, to tackle (C1), we utilize the Mixture-of-Experts (MoE) framework (Bian et al., 2023) to integrate semantic-guided knowledge experts, constructing expert networks for entities across various relational contexts. Additionally, we implement a joint interaction framework that leverages appropriate contexts and captures the intricate connections between entity-relation pairs and attributes within a unified vector space. Employing this knowledge-aware mechanism, our model discerns the significance and relevance of each attribute, effectively synthesizing the latent semantics among them. To address (C2), diverging from prior work that primarily focuses on positive and negative samples, we introduce a novel ordinal knowledge contrastive learning method to generate high-quality ordinal samples. We optimize these ordinal samples based on cosine similarity, selecting the top-k samples with the highest similarity to enhance the learning of ordinal relations. This approach enables the model to capture the continuous features of numerical attributes and discern the subtle differences in boundary values, thus enhancing numerical reasoning. Furthermore, we provide theoretical guarantees for the effectiveness of our model. Extensive experiments show our model can achieve the state-of-the -art performance.
In summary, our contributions are three-fold:
  • To conduct knowledge-aware attribute learning, we introduce a novel approach, KAAE, from semantic interaction and perceptual perspectives. To our knowledge, this is the first model to address semantic relevance by integrating entity relations and numerical attributes within a unified framework for numerical reasoning tasks.
  • To capture the fine-grained semantics of continuous features for numerical attributes, we propose a novel ordinal knowledge contrastive learning method that generates high-quality ordinal samples. This approach enhances the model’s ability to perform nuanced numerical reasoning by exploiting ordinal relationships within the data.
  • Extensive experiments show that KAAE achieves state-of-the-art performance on three real-world datasets, and demonstrate its robustness, reasonability, and interpretability.
2. Related Work
Knowledge Graph Embedding(KGE). KGE amis to capture the latent representations of entities and relations in KGs. TransE(Bordes et al., 2013) and its extended models(Wang et al., 2014; Zhang et al., 2021; Ji et al., 2015), focus on treating relation as a ”translation” from the head entity to the tail entity. DistMult(Dong et al., 2014) represent each relation with a diagonal matrix, ComplEx(Trouillon et al., 2016) extends DistMult to complex space. ConvE(Dettmers et al., 2018) extracts deep features of head entity and relation based on 2D convolution. ConE(Bai et al., 2021), MuRP(Balazevic et al., 2019a), GIE(Cao et al., 2022b) and AttH(Chami et al., 2020) embed KGs into hyperbolic spaces to model hierarchical relations. Further, models(Schlichtkrull et al., 2018; Shang et al., 2019; Vashishth et al., 2020) of GNNs have been proposed to model higher-order connectivity in knowledge graphs. However, these models do not consider numerical values, making it difficult to accomplish numerical reasoning tasks.
Knowledge Graph Embedding with Numerical Attributes. To improve the performance of KGE, several KGE models have been proposed to incorporate auxiliary information about the entity such as literals and numerical attributes (e.g., age, weight, and born_year). For example, KBLRN(García-Durán and Niepert, 2018) accomplish the KGE task based on numerical values, considering numerical differences between different entities. MT-KGNN(Tay et al., 2017) and TransEA(Wu and Wang, 2018) are models for multi-task learning to predict both numeric attribute values and entity/relation embeddings. LiteralE(Kristiadi et al., 2019) combines the entity embedding with numeric attributes based on a learnable gating function. Deterministic stand-alone value representation methods including DICE(Sundararaman et al., 2020), NEKG(Duan et al., 2021) and NRN(Bai et al., 2023) are also used to predict numerical attributes in KGs. Moreover, RAKGE(Kim et al., 2023) learns the representation for attribution with relations. However, these methods do not consider the joint information among entities, attributes, and relations. On the contrary, our KAAE model can exploit joint complex interactions among entities, relations and numerical attributes to fully mine the semantic information useful for numerical reasoning tasks.
Contrastive Learning. Contrastive learning (CL) is an effective self-supervised learning method by pulling semantically close neighbors together while pushing non-neighbors away, and some models have attempted to utilize contrastive learning to improve representations of entities/graphs. SimGCL(Yu et al., 2022) adds uniform noise to the entity representation, which is an expansion-free CL method. HeCO(Wang et al., 2021b) employs contrast learning in network schema and meta-path, capturing both local and global features of entities. SLiCE(Wang et al., 2021a) makes use of mutual attraction between closed nodes to learn subgraph representations. RAKGE(Kim et al., 2023) interacts with training samples with head entities to generate samples. However, none of the above methods can guarantee effectiveness in numerical reasoning tasks. To improve contrast learning, we optimize the sampling process with ordinal relations in KGs and generate high-quality ordinal positive/negative samples based on the technique of similarity sampling, thus achieving significant performance improvements.
3. Methodology
In this section, we present a new KG-based numerical reasoning framework called KAAE from the semantic joint interaction perspective. Specifically, there are two main components: Mixture-of-Experts-Knowledge-Aware Encoder (MoEKA-Encoder) and ordinal knowledge contrastive learning (OKCL) strategy. The overall framework of KAAE is illustrated in Figure 2.
Problem Definition. Given a knowledge graph as a collection of real triples $\mathcal{G}=\{(h,r,t)\}$, where $h,t\in\mathcal{E}$ and $r\in\mathcal{R}$ represent the set of entities and relations, respectively. Denote an entity-numeric value matrix for entities as $\bm{X}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{M}|}$ and the m-th numeric value belongs to the entity i as $X_{im}$, where $\mathcal{M}$ is the set of numeric attribute fields (e.g., height, weight, and born_year). Due to the incompleteness of the KG, KGE emerge as an essential approach to conduct the knowledge graph completion, which is to map each triple (h, r, t) to a reasonableness score based on the KGE model, where a low score means that the triple is not reasonable, and vice versa. In this paper, our purpose is to construct a new KGE model to conduct numerical reasoning with the assistance of numerical attributions.
Input Knowledge Graph             MoEKA Encoder          Ordinal Knowledge Contrastive Loss
Figure 2. Overview of our KAAE model. The MoEKA Encoder encodes each entity with the relation and attributes to a unified representation thus tackling the semantic relevance. Afterward, to capture the fine-grained semantic information, we utilize a new knowledge contrastive learning method to generate high-quality ordinal samples thus tackling semantic ambiguity. In this way, we can learn the nuanced differences in attributes and perform better numerical reasoning tasks.
3.1. Knowledge-Aware Learning
To tackle the semantic relevance of KGs, we propose a MoE-knowledge-aware encoder consisting of four parts: the numeric value embedding layer, a mixture of knowledge experts, the knowledge perceptual attention layer, and the knowledge aggregation layer.
Numeric Value Embedding Learning. To learn the numeric values, for observable scalars, we use a learnable embedding matrix to map them to a vector space. For missing scalars, we use learnable special missing value embeddings to preserve the magnitude information and prevent the task from being threatened, instead of using fixed values such as zero: $\bm{o}_{i}^{m}=(\bm{W}_{x}^{m}+\bm{w}_{m}X_{im})\odot\bm{v}_{m}$, where $X_{im}$ is the corresponding numeric value of the entity i. $\bm{v}_{m}\in\mathbb{R}^{d_{att}}$ represents the embedding vector of the numeric attribute field m, $d_{att}$ stands for the embedding dimension of attribute. $\bm{w}_{m}\in\mathbb{R}^{d_{att}}$ and $\bm{W}_{x}^{m}\in\mathbb{R}^{d_{att}\times d_{att}}$ are linear transformations that learn the context information between each numeric value and attribute. The operation symbol $\odot$ represents point-wise multiplication. We denote $\bm{o}_{i}^{m}\in\mathbb{R}^{d_{att}}$ as the m-th field attribute embedding of the entity i. In this way, we can learn a more robust representation of both existing/missing numerical attributes.
Mixture of Knowledge Experts. To better learn entity embeddings from different perspectives with relation contexts and numerical attributes, we first employ a module called Mixture of Knowledge Experts to build expert networks, where each perspective corresponds to an expert network. First, the entity $i\in\mathcal{E}$ has a raw feature $\tilde{e}_{i}$. We then learn the multi-perspective embeddings $\mathcal{H}^{\tilde{e}_{i}}_{1},\mathcal{H}^{\tilde{e}_{i}}_{2},\ldots,\mathcal{H}^{\tilde{e}_{i}}_{K}$ for the entity $i$ by establishing $K$ knowledge experts denoted as $\mathcal{W}_{1},\mathcal{W}_{2},\ldots,\mathcal{W}_{K}$. This process can be represented as $\mathcal{H}^{\tilde{e}_{i}}_{k}=\mathcal{W}_{k}(\tilde{e}_{i})$. Next, we design a semantic-guided gated fusion network (SGN) to facilitate the fusion of inter-perspective entity embeddings with relation guidance.
(1) $\bm{e}_{i}=\sum\nolimits_{k=1}^{K}\mathcal{F}_{k}(\mathcal{H}^{\tilde{e}_{i}}_{k},r)\mathcal{H}^{\tilde{e}_{i}}_{k}$
where $\bm{e_{i}}\in\mathbb{R}^{d_{emb}}$ is the output entity embedding of entity i for relation $r$, $\mathcal{F}_{k}$ is the weight for each expert calculated by the GFN:
(2) $\mathcal{F}_{k}(\mathcal{H}^{\tilde{e}_{i}}_{k},r)=\frac{exp((\mathcal{U}(\mathcal{H}^{\tilde{e}_{i}}_{k})+\psi_{k})/\rho(\epsilon_{r}))}{\sum\nolimits_{j=1}^{K}exp((\mathcal{U}(\mathcal{H}^{\tilde{e}_{i}}_{j})+\psi_{j})/\rho(\epsilon_{r}))},$
where $\psi_{k}\sim\mathcal{N}(0,\mathcal{U^{\prime}}(\mathcal{H}^{\tilde{e}_{i}}_{k}))$, $\mathcal{U}$ and $\mathcal{U^{\prime}}$ are two projection layers, and $\psi_{k}$ is tunable Gaussian noise used to balance the weights for each expert and enhance the model’s robustness. Additionally, we add a relation-aware temperature $\epsilon_{r}$ with a sigmoid function $\rho$ to limit the temperature within the range $(0,1)$. Our aim is to obtain an entity embedding within the relational context of the current prediction before making the final decision. This approach introduces the relational context to guide embedding learning within the experts, enabling them to extract relation-aware embeddings. Additionally, the experts will be differentiated to adapt to different relational contexts through the design of SGN. This allows us to learn dynamic embeddings of entities that change according to different relational contexts.
Knowledge Perceptual Attention. In the previous sections, we obtained the entity representations within the relational context. Based on this, we aim to learn the knowledge-aware representations via the joint semantic interactions among entities, relations, and attributes. In this way, we propose a knowledge perceptual attention mechanism for entity-relation attributes using multi-head self-attention(Vaswani et al., 2017) to capture the importance of numerical attributes based on each entity and its associated relations. Specifically, we leverage a single layer perception $f_{e-att}$ and $f_{r-att}$ to project the entity and relation embeddings onto the attribute embedding space, respectively. Thereafter, we mixing the information of entity e and relation r, so as to get the joint projected embedding $\bm{p}_{joint}^{att}\in\mathbb{R}^{d_{att}}$:
(3) $\bm{p}_{joint}^{att}=\sum_{*\in\{e,r\}}\delta_{*}f_{*-att}(\bm{p}_{*})$
where $\bm{p}_{e}$ corresponds to $\bm{e}_{i}$, $\bm{p}_{r}$ corresponds to the relation embedding, $\delta_{*}\in[0,1]$ are hyperparameters and $\delta_{e}+\delta_{r}=1$.
Then, we apply the knowledge perceptual attention mechanism to establish the joint interactions between entities, relations, and attributes. Specifically, We apply this mechanism for all attribute fields $m$ = $\cdots,|\mathcal{M}|$ and we repeat this step for all attention heads l = 1,2$\cdots$,$L$, where $L$ is the number of multi-head attentions. The result of the joint interaction of head l is shown as follows:
(4) $\bm{o}_{joint,i}^{(l)}=\sum\nolimits_{m=1}^{|\mathcal{M}|}a_{joint,i,m}^{(l)}(\bm{W}_{agg}^{(l)}\bm{o}_{i}^{m}),$
where the linear transformation matrix $\bm{W}_{agg}^{(l)}\in\mathbb{R}^{d_{sub}\times d_{att}}$ aims to project attribute embedding $\bm{o}_{i}^{m}$ into low-dimensional subspaces to capture the importance of each numeric attribute from the given entity with relation accurately. Under a specific attention head l, we map $\bm{e}_{joint}^{att}$ and $\bm{o}_{i}^{1},\cdots\bm{o}_{i}^{|\mathcal{M}|}\in\mathbb{R}^{d_{att}}$ onto smaller subspaces and capture the attention score. The normalized attention weight $a_{joint,i,m}^{(l)}$ of the attribute embedding $\bm{o}_{i}^{m}$ is formulated as follows:
(5) $a_{joint,i,m}^{(l)}=\frac{exp(s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m}))}{\sum\nolimits_{n=1}^{|\mathcal{M}|}exp(s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{n}))},$
and
(6) $s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m})=\sigma\left(\|_{j=1}^{H}\bm{W}_{j}^{(l)}\phi_{j}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m})\right)$
[…]
(7) $\bm{e}_{joint,i}^{att}=\sigma(\bm{W}_{ge}\bm{e}_{i}+\bm{W}_{ga}\bm{o}_{joint,i})+\bm{b}$
where $\bm{W}{ge}\in\mathbb{R}^{d_{emb}\times d_{emb}}$ and $\bm{W}_{ga}\in\mathbb{R}^{d_{emb}\times d_{att}}$ are different linear transformations to assign different weights to each entity representation and fused numerical representation by adaptively learning to obtain high quality final embeddings and to focus on the more important information. $\bm{b}$ is a bias. $\sigma$ is a sigmoid function, and $\tau$ is a hyperbolic tangent function. We name $\bm{e}_{joint,i}^{att}\in\mathbb{R}^{d_{emb}}$ as the attribute-enriched vector of entity i.
3.2. Ordinal Knowledge Contrastive Learning
Learning ordinal relations is essential for numerical reasoning tasks. Specifically, for numerical reasoning, data consists of continuous scalar values, resulting in numerical values that are extremely sensitive to minimal differences near boundary values, thus affecting the accuracy of ordinal embedding. Unlike the previous work mainly focusing on generating valid, positive, and negative samples, we turn to a more difficult and significant task that generates high-quality ordinal samples. For example, consider the head entity with a weight attribution of 80kg and the relation is_heavier_than. The ideal learning range consists of generating all possible positive tail entities weighing less than 80kg. It is particularly crucial to distinguish boundary values(e.g., tails between 79 and 81kg) more efficiently.
Ordinal Relation Learning. Given $N$ objects and the ordinal relationship triples $\mathcal{S}=\{(a,b,c)\mid a,b,c\in\mathcal{E},a\neq b\neq c\}$, where $[N]=\{1,\ldots,N\}$ and object $a$ is more similar to object $b$ than it is to object $c$. Given some distance function $d(\cdot,\cdot)$, we aim to learn the representations of objects, denoted as $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$, such that the following objectives hold as much as possible:
(8) $d(\mathbf{x}_{a},\mathbf{x}_{b})<d(\mathbf{x}_{a},\mathbf{x}_{c}),\quad\forall(a,b,c)\in\mathcal{S}.$
Knowledge Samples Generator. On the analysis above, to distinguish precise semantics in numerical reasoning, we generate the preliminary positive/negative ($\mathcal{E}^{+}$ and $\mathcal{E}^{-}$) samples based on the available head entities and relations, the set of positive/negative samples is as follows:
(9) +=$\centering\mathcal{E^{+}}=\{\bm{e}_{joint,i}^{att}|i\in\mathcal{P}[h,r]\},\ \centering\mathcal{E^{-}}=\{\bm{e}_{joint,j}^{att}|j\in\mathcal{N}[h,r]\},\@add@centering\@add@centering$
where $\mathcal{P}[h,r]$ represents the set of positive tails containing entities whose weight is less than 80kg (e.g., 65, 70, and 75kg). $\mathcal{N}[h,r]$ represents the set of positive tails containing entities whose weight is more than 80kg (e.g., 85, 90, and 95kg).
Ordinal Samples Extractor. Since each attribute field (e.g., year, age, weight) has its unique distribution, we select k number of samples with the highest cosine-similarity to $\bm{e}_{joint,i}^{att}$ from all the generated positive/negative samples with the assistance of Eq.(8). Note that this step is essential and it can generate the samples obeying ordinal embedding. Then we generate high-quality ordinal samples as follows:
(10) $\bm{e}_{mix}^{t+}=\alpha\cdot\sum_{p\in Topk(\mathcal{E^{+}})}\bm{e}_{joint,p}^{att}+(1-\alpha)\cdot\bm{e}_{joint,h}^{att},$
(11) $\bm{e}_{mix}^{t-}=\beta\cdot\sum_{n\in Topk(\mathcal{E^{-}})}\bm{e}_{joint,n}^{att}+(1-\beta)\cdot\bm{e}_{joint,h}^{att},$
where $\alpha,\beta$ are blending coefficients that are sampled from the uniform distribution $[0,1]$. $Topk(\mathcal{E^{+}})$ and $Topk(\mathcal{E^{-}})$ are sampling sets with the highest cosine similarity to $\bm{e}_{joint,i}^{att}$ in $\mathcal{E}^{+}$ and $\mathcal{E}^{-}$, respectively. In this sense, to satisfy ordinal relation, we have:
$d(\bm{e}_{joint,h}^{att},p)<d(\bm{e}_{joint,h}^{att},n)$
where $p\in Topk(\mathcal{E^{+}})$ and $n\in Topk(\mathcal{E^{-}})$. In this way, our contrastive learning loss is defined as follows:
(12) $\begin{split}\mathcal{L}_{CL}=-&\frac{1}{\mathcal{|G|}}\sum\nolimits_{l\in\mathcal{G}}\log\frac{exp(\textbf{p}^{\top}\textbf{u}/\tau)}{exp(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\bm{e}_{mix}^{t-}}exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)},\end{split}$
where $\bm{p}=\bm{e}_{joint,h}^{att}+\bm{e}_{r}$, $\bm{u}=\bm{e}_{mix}^{t+}$, and $\tau$ is the temperature hyper-parameter. By implementing the strategy of randomly mixing the positive tails and employing head blending techniques (as described in Eqs. (10)-(11)), we notably increase the diversity of our samples. Consequently, this strategy significantly refines our model’s ability to navigate and interpret the nuances of semantic information, thereby reducing semantic ambiguity and bringing a more accurate and profound comprehension of semantics, advancing beyond the limitations imposed by the initial training data.
3.3. Training
Score Function. A classical assumption of typical KGE methods is to compute the distances between head entities, relations, and tail entities. However, in numerical reasoning tasks, these methods suffer from serious drawbacks because they do not effectively address the order embedding problem posed by relational patterns in complex tasks. For example, given the head entity (weight=80kg) and its relation (is_heavier_than), and using the TransE score function, the two tail entities $e_{1}$ (weight=70kg) and $e_{2}$ (weight=60kg) are mapped to the same location, failing to express a numerical ordering relation such as 80>70>60, which impedes numerical reasoning. To address this issue, we design a rationality score function specifically for the numerical reasoning task as follows:
(13) $\begin{split}score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{joint,i}^{att})=&\varepsilon-\|\bm{e}_{joint,h}^{att}+\bm{e}_{r}-\bm{e}_{joint,i}^{att}\|_{1/2}+\\ &\Delta\|max(0,\bm{W}_{r}\bm{e}_{joint,h}-\bm{W}_{r}\bm{e}_{joint,i}^{att})\|^{2},\end{split}$
where $\Delta$ represents the weight in the numerical reasoning score function. The first term is derived from TransE (Eq. (17)), while the second term modifies the Order-embedding approach (Eq. (18)). $\bm{W}_{r}\in\mathbb{R}^{d_{emb}\times d_{emb}}$ is a projection matrix critical for mapping entities to a specific relational space, indicating that depending on the type of relation, each entity may assume a different order.
Loss Function. In this paper, we use the binary cross-entropy loss that presented in ConvE(Dettmers et al., 2018) and LiteralE(Kristiadi et al., 2019) so that we can convert the score of each triple obtained based on Eq.(13) to a probability by using a sigmoid function.Let $\mathcal{T}=\mathcal{G}\cup\mathcal{G}^{-}$ denote the training dataset,where $\mathcal{G}$ denotes the set of positive knowledge triples, $\mathcal{G}^{-}$ denotes the set of negative knowledge triples $\{(h,r,t^{{}^{\prime}})|h,t^{{}^{\prime}}\in\mathcal{E},r\in\mathcal{R},(h,r,t^{{}^{\prime}})\notin\mathcal{G}\}$. The binary-cross entropy loss is defined as follows:
(14) $\mathcal{L}_{BCE}=-\frac{1}{|\mathcal{T}|}\sum\limits_{l\in\mathcal{T}}(y_{l}\log(p_{l})+(1-y_{l})\log(1-p_{l})),$
where $y_{l}\in\{0,1\}$ is the truth label, and $p_{l}\in$ [0,1] is the probability of each triple $(h,r,t)=l\in\mathcal{T}$, which is formulated as Eq.(13). Combining Eq. (12) and Eq. (14), the final loss can be summarized as follows:
(15) $\mathcal{L}_{total}=\mathcal{L}_{BCE}+\lambda\mathcal{L}_{CL},$
where $\lambda$ is a hyperparameter, stands for the coefficient of contrastive learning loss.
Table 1. ‘Numeric-aware’ indicates the availability of numeric information, while ‘Relation-aware’ indicates learning the presence of relation-specific. Knowledge-aware indicates Interacting with entities and relations simultaneously.
TransE BiGI R-GCN LiteralE RAKGE KAAE
[…]
Table 2. Results for numerical reasoning. Bold scores indicate the best results, while underlined scores represent the second-best results. The % of Improvement column shows the relative improvements of KAAE compared to the second-best scores.
Model US-Cities Spotify Credit
H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR
Euclidean TransE 0.189 0.248 0.324 367 0.239 0.259 0.355 0.462 115 0.332 0.421 0.520 0.630 39 0.493
DistMult 0.013 0.028 0.065 855 0.032 0.025 0.048 0.131 227 0.062 0.127 0.195 0.301 122 0.189
ComplEx 0.084 0.126 0.185 587 0.120 0.124 0.188 0.284 173 0.182 0.291 0.388 0.513 65 0.366
ConvE 0.158 0.202 0.274 385 0.198 0.231 0.307 0.414 110 0.295 0.171 0.281 0.430 58 0.261
TuckER 0.156 0.212 0.308 321 0.207 0.211 0.296 0.411 98 0.278 0.405 0.512 0.638 36 0.485
HAKE 0.003 0.025 0.064 981 0.024 0.008 0.085 0.120 196 0.075 0.051 0.151 0.273 128 0.134
Hyperbolic AttH 0.051 0.069 0.108 1255 0.076 0.052 0.082 0.146 348 0.087 0.176 0.261 0.395 101 0.251
ConE 0.009 0.042 0.114 239 0.048 0.001 0.006 0.064 227 0.028 0.006 0.241 0.395 82 0.156
GIE 0.095 0.134 0.206 570 0.132 0.121 0.183 0.281 185 0.174 0.285 0.388 0.512 76 0.359
MuRP 0.082 0.113 0.175 457 0.115 0.024 0.179 0.324 135 0.119 0.151 0.246 0.432 80 0.228
GNNs R-GCN 0.212 0.271 0.355 314 0.263 0.288 0.382 0.504 89 0.364 0.480 0.570 0.675 34 0.546
WGCN 0.029 0.056 0.104 968 0.057 0.095 0.162 0.261 331 0.153 0.170 0.256 0.369 100 0.241
Attributed KBLRN 0.006 0.018 0.046 2164 0.021 0.017 0.039 0.086 347 0.044 0.007 0.020 0.079 271 0.061
MT-KGNN 0.071 0.109 0.156 653 0.102 0.108 0.182 0.302 142 0.176 0.211 0.304 0.436 74 0.285
LiteralE 0.246 0.308 0.402 228 0.299 0.264 0.371 0.498 76 0.345 0.475 0.562 0.676 36 0.549
Deterministic DICE 0.182 0.224 0.284 368 0.209 0.239 0.312 0.412 126 0.320 0.402 0.509 0.597 38 0.482
NEKG 0.196 0.230 0.304 376 0.217 0.246 0.323 0.423 121 0.336 0.398 0.512 0.589 41 0.477
NRN 0.298 0.326 0.432 155 0.323 0.342 0.442 0.546 102 0.414 0.560 0.639 0.740 21 0.620
Self-supervised BiGI 0.185 0.249 0.331 359 0.236 0.260 0.354 0.468 118 0.331 0.418 0.507 0.622 39 0.487
SliCE 0.185 0.250 0.331 359 0.237 0.261 0.354 0.469 117 0.332 0.420 0.510 0.622 38 0.490
SimGCL 0.344 0.415 0.502 162 0.399 0.000 0.255 0.467 59 0.167 0.000 0.399 0.645 22 0.239
RAKGE 0.395 0.455 0.529 199 0.442 0.502 0.608 0.674 37 0.573 0.647 0.733 0.823 12 0.708
KAAE 0.439 0.536 0.640 119 0.508 0.636 0.694 0.758 30 0.680 0.745 0.818 0.888 7 0.794
lmprovement 11.1% 17.8% 21.0% 26.5% 14.9% 26.7% 14.1% 12.5% 18.9% 18.7% 15.1% 11.6% 7.9% 41.7% 12.1%
3.4. Theoretical Analysis
To clarify how our model differs from these models, we compare our model with some popular models in Table1.
Time Complexity Analysis. KAAE consists of three components, the MoEKA encoder, a process of contrastive learning, and the score function. In terms of MoEKA Encoder, the time complexity of processing every triple is $\mathcal{O}(|\mathcal{M}|)$. The time complexity of the numerical value embedding layer is $\mathcal{O}(|\mathcal{M}|)$ because we transform each attribute field into an embedding vector. In the joint perceptual attention layer, the query is the target relation and the number of keys and values corresponds to the number of attribute fields, so the time complexity remains $\mathcal{O}(|\mathcal{M}|)$. The gating layer of the MoEKA Encoder is not involved in the calculation of time complexity. During the process of contrastive learning, for every given triple(h,r,t), the time complexity is proportional to the number of positive samples $|\mathcal{P}[h,r]|$ and the number of negative samples $|\mathcal{N}[h,r]|$. The time complexity is $\mathcal{O}(|\mathcal{M}|\cdot rel\_{}num)$, where rel_num denotes the number of relations per entity. The TransE score function involves only addition and subtraction operations, so it does not change the time complexity. Therefore, the overall time complexity of KAAE is $\mathcal{O}(\mathcal{T}\cdot(|\mathcal{M}|+|\mathcal{M}|\cdot rel\_{}num))\approx\mathcal{O}(|\mathcal{T}|\cdot rel\_{}num)$, where $\mathcal{T}$ denotes the number of positive and negative triples used for training.
The Expressiveness of Our Model. We conduct the expressiveness of numerical reasoning from the perspective of a one-hop reasoning task. Specifically, we aim to theoretically analyze the process of contrastive learning in KAAE and establish theoretical guarantees for the downstream performance of the learned representations. In the numerical reasoning, we denote the set of node representations as $\mathcal{V}$, and define the mean representations from one-hop neighborhoods of node $v$ as $\bm{z}=\frac{1}{\mathcal{N}(v)}\sum\nolimits_{\bm{u}\in\mathcal{N}(v)}\bm{u}$, where $\bm{u}$ represents the projected node representations of positive samples of node $v$, $\bm{z}$ describes the one-hop neighborhood pattern of node $v$. In the world, nodes belonging to the same semantic class tend to have similar neighborhood patterns, so $\bm{z}$ can be viewed sampled from $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$ where $Y$ is the latent semantic class indicating the one-hop pattern of node $v$. We demonstrate that minimizing KAAE’s objective in Equation (12) with an exponential moving average is equivalent to maximizing mutual information between representation $V$ and the one-hop pattern $Y$, which explains the rationality of KAAE in capturing one-hop patterns as follows:
(16) $\begin{split}\mathcal{L}_{CL}\geq H(V|Y)-H(V)=-I(V;Y),\end{split}$
Eq. (16) indicates minimizing KAAE loss in Eq. (12) promotes maximizing the mutual information $I(V;Y)$ between representations and one-hop neighborhood context. In this way, our KAAE can exploit more latent semantics among KGs, hence facilitating numerical reasoning. All the detailed proofs are in Appendix B.
4. Experiments
4.1. Experimental Settings
Datasets. We use three real-world KG datasets to evaluate the proposed model on a numerical reasoning task. Specifically, US-Cities [^1] https://simplemaps.com/data/us-cities is a large-scale city dataset containing basic information about several cities in the United States, including more than 70 attribute fields (e.g., household income, household education level, dimension,etc). Spotify is a knowledge graph provided by Spotify for developers [^2] https://www.kaggle.com/datasets/geomack/spotifyclassification. Credit is a knowledge graph constructed from default events in Taiwan(Yeh and Lien, 2009). Each credit card holder corresponds to an entity with numerical information about the credit card (e.g., total_delay, total_bill) as the numerical attribute. Additionally, it is worth noting that all three of the above datasets are modified to simulate numerical reasoning tasks under real conditions. Close to 20% of the numerical values are masked to zero (missing values) while using as few triples as possible for the training set. The remaining triples are used for evaluation. The statistics of the dataset can be found in Appendix C.
Evaluation Metrics. We used the improved learning methods based on dropout, batch normalization, and linear transformation proposed alongside LTE(Zhang et al., 2022) for evaluation. Mean Reciprocal Rank (MRR), Mean Rank (MR) and Hits@{1, 3, 10} are reported as the performance metrics to evaluate the proposed model’s performance. Higher Hit@n and MRR values indicate a better performance, whereas lower MR values imply a better performance.
Baseline. We compare KAAE with six groups of existing representation learning methods: Euclidean KGE including TransE(Bordes et al., 2013), ConvE(Dettmers et al., 2018), ComplEx (Trouillon et al., 2016), TuckER(Balazevic et al., 2019b), DistMult(Dong et al., 2014) and HAKE(Zhang et al., 2020). Hyperbolic KGE including MuRP(Balazevic et al., 2019a), ConE(Bai et al., 2021), AttH(Chami et al., 2020) and GIE(Cao et al., 2022b); GNNs-KGE including R-GCN(Schlichtkrull et al., 2018) and WGCN (Shang et al., 2019); Attributed KGE including KBLRN(García-Durán and Niepert, 2018), MT-KGNN (Tay et al., 2017) and LiteralE(Kristiadi et al., 2019); and Self-supervised model for graph including BiGI(Cao et al., 2021), SLiCE(Wang et al., 2021a), SimGCL(Yu et al., 2022), and RAKGE(Kim et al., 2023); Deterministic stand-alone value representation methods including DICE(Sundararaman et al., 2020), NEKG(Duan et al., 2021) and NRN(Bai et al., 2023). Among all the baselines, for those that do not take care of external numerical attributes, such as TransE, we follow their original approach and concat numerical values to the entity embeddings. The results of Self-supervised KGE are from RAKGE(Kim et al., 2023). More details for implements can refer to the Appendix C.
4.2. Experimental Results
Performance Comparison. To address these challenges, we introduce a new model termed KAAE (Knowledge-Aware Attributes Embedding), crafted for numerical reasoning via knowledge graphs. Specifically, to tackle semantic relevance, we utilize the Mixture-of-Experts (MoE) framework (Bian et al., 2023) to integrate semantic-guided knowledge experts, constructing expert networks for entities across various relational contexts. Additionally, we implement a joint interaction framework that leverages appropriate contexts and captures the intricate connections between entity-relation pairs and attributes within a unified vector space. Employing this knowledge-aware mechanism, our model discerns the significance and relevance of each attribute, effectively synthesizing the latent semantics among them. To address semantic ambiguity, diverging from prior work that primarily focuses on positive and negative samples, we introduce a novel ordinal knowledge contrastive learning method to generate high-quality ordinal samples. We optimize these ordinal samples based on cosine similarity, selecting the top-k samples with the highest similarity to enhance the learning of ordinal relations. This approach enables the model to capture the continuous features of numerical attributes and discern the subtle differences in boundary values, thus enhancing numerical reasoning. Furthermore, we provide theoretical guarantees for the effectiveness of our model. Extensive experiments substantiate the effectiveness of our model.
To validate the effectiveness of our ordinal knowledge contrastive learning method, we compared KAAE with those self-supervised models. MRR is improved from 0.331 for BiGI and 0.573 for RAKGE to 0.680 on Spotify dataset, we can deduce KAAE learn richer representations than graph-based approaches including BiGI, SLiCE, and SimGCL by considering extra numerical attributes of entities. Compared to RAKGE, KAAE focuses on generating and extracting hard high-quality ordinal samples to distinguish boundary numerical values and magnitudes accurately, which makes contrastive learning more effective in numerical reasoning tasks.
Besides, while HAKE, TuckER, and ConvE often outperform TransE and DistMutl models, Table 2 suggests otherwise. We speculate that this is because numerical reasoning is a one-to-one task in most cases, and is better suited to TransE and DistMult.
4.3. Experimental Analysis
Ablation Study. We conduct an ablation experiment to evaluate the performance of different variants of KAAE. Initially, we set $\delta_{e}$ in Eq.(3) to 0 to disable the entity-aware module. This modification allows the model to perceive numerical attributes solely through a relation-aware approach, without considering the embeddings of entities associated with the attributes. We also test the performance of KAAE without a full understanding of the entity semantics through the MoE module. Furthermore, we train the model without incorporating the top-k ordinal sampling methodology (Eqs. (10-11)) into the contrastive learning framework, which is essential for generating high-quality training samples and enhancing the model’s ability to discern discrete numerical boundaries. The results, as detailed in Table 3, indicate that the omission of any single module results in a decline in model performance. This finding emphasizes the critical role of making the most of entity-related information in numerical reasoning tasks and the significance of high-quality ordinal samples in contrastive learning.
Table 3. Ablation study of KAAE on Credit dataset. ‘MoE’ stands for Mixture-of-Experts, ‘KA’ stands for Knowledge-Aware, ‘OS’ stands for Ordinal-Sampling.
MoE KA OS H@1 H@3 H@10 MR MRR
[…]
Table 4. Case study on Credit dataset. IDXX means that we consider the information from an entity with a definite relation
Model Relation Entity Important Attributes(Attention Scores)
KAAE TOTAL_BILL_comp ID1763 TOTAL_BILL(0.278), LIMIT_BAL(0.243), AGE(0.193), TOTAL_DELAY(0.160), TOTAL_PAY(0.127)
MARRIAGE ID551 AGE(0.327), TOTAL_BILL(0.235), TOTAL_DELAY(0.207), TOTAL_PAY(0.144), LIMIT_BAL(0.088)
EDUCATION ID1130 AGE(0.356), TOTAL_DELAY(0.232), TOTAL_BILL(0.192), TOTAL_PAY(0.126), LIMIT_BAL(0.094)
RAKGE TOTAL_BILL_comp - AGE(0.203), TOTAL_PAY(0.202), TOTAL_BILL(0.201), LIMIT_BAL(0.197), TOTAL_DELAY(0.196)
MARRIAGE - TOTAL_PAY(0.202), TOTAL_BILL(0.2019), AGE(0.2017), TOTAL_DELAY(0.1971), LIMIT_BAL(0.1970)
EDUCATION - LIMIT_BAL(0.201), TOTAL_DELAY(0.200), TOTAL_PAY(0.1998), TOTAL_BILL(0.1997), AGE(0.1996)
Analysis of MoEKA Encoder. As mentioned above, the crucial role of our MoEKA Encoder is to extract the numerical information contained in the entity and relation with attributes, to accomplish joint perception with each attribute. The attention matrix, obtained using Eq. (5), is visualized in Figure 3. To make it easier to observe, we select a portion of all entities to visualize the attention scores, the first heatmap shows the entity-attribute association scores (setting $\delta_{r}$ in Eq. (3) to 0). The second heatmap shows the association scores after adding information from a relation to each entity (setting $\delta_{e}$ and $\delta_{r}$ to 0.5) in the first heatmap, effectively improving the problem of equalizing the attention to attributes for some entities. For example, Entity ID156 has close association scores with all 4 attributes except TOTAL_PAY, and taking into account the information of the relation TOTAL_PAY_comp, the MoEKA Encoder can better differentiate between these 5 attributes. In addition, we did another similar experiment to observe the effect of entity information on relation-attribute attention scores. The results are shown in Figure 7 in the Appendix, indicating that adding information from an entity to each relation effectively improves the problem of equalizing the attention to attributes in some relations.
We also investigate the effect of the number of experts K in the MoE module, as depicted in Figure 4. It can be observed that the impact of the number of experts K on the final results generally follows a pattern of initial increase followed by a decrease, mainly affecting fine-grained metrics such as Hit@1 and MRR. Having either too many or too few experts is detrimental to the model’s learning performance. The model achieves the best results at K = 3.
Figure 3. Visualization of relevance scores of each numeric attribute without/with relation-aware on Credit dataset.
Figure 4. Research of Hyperparameter K on Spotify dataset.
Case Study. The MoEKA Encoder in KAAE was designed to obtain more robust attribute representations using the relevant context. Therefore, we investigated significant attributes contributed to the inference of a derived relation. Table 4 presents the attention scores of the MoEKA Encoder for all the attributes used to infer the comparison relations in the Credit dataset. When predicting relation TOTAL_BILL_comp, KAAE was able to focus more attention on the TOTAL_BILL attribute associated with TOTAL_BILL_comp. However, RAKGE(Kim et al., 2023) did not do so, paying essentially the same amount of attention to different attributes, which were all around 20%. Moreover, age is clearly a very important factor when predicting relations MARRIAGE and EDUCATION. By taking into account the information from entity embeddings, KAAE was able to focus more on the AGE attribute, while previous model RAKGE can not.
Figure 5. Proportional test of KAAE on the Credit dataset. A relation percentage of 100% means that we only consider relations, while 0% means that we only consider entities.
Proportional Analysis. Based on Eq. (3), we alter the proportion of weights between entities and relations in joint perception, exploring the improvement of numerical reasoning by the numerical information contained in entities and relations. Figure 5 illustrates the performance of KAAE when relations occupy different weights in joint perception, and the results show that the model achieves excellent performance when the proportion of relations is 40%-90%. There is an improvement of 46.1% in MRR and 60.2% in Hits@1 when the proportion of relations is 90% than when only entities are considered. However, the model performance deteriorates when we only consider relations and do not focus on entities. We can infer that both entities and relations contain information that is helpful for numerical reasoning. We need joint interaction of entity-relation attributes to accomplish numerical reasoning tasks efficiently.
Extreme value Analysis. To test whether our edge-range sampling approach will lead to a weaker perception of large numerical differences, we did an additional experiment to validate KAAE by selecting different proportions of positive/negative samples separately on Credit dataset. Table 5 demonstrates that KAAE can still maintain a good performance without utilizing all samples.
Table 5. Performance with different proportions of samples.
Proportion H@1 H@3 H@10 MR MRR
20% 0.739 0.805 0.882 7 0.796
40% 0.737 0.802 0.879 7 0.782
60% 0.746 0.821 0.885 7 0.789
80% 0.743 0.812 0.876 7 0.792
100% 0.745 0.818 0.888 7 0.794
Additionally, we added 10% extra extreme samples to the datasets. The results in Table 6 show that extreme values will not affect the performance of KAAE.
Table 6. Performance of KAAE with extra extreme values.
Dataset Extreme H@1 H@3 H@10 MR MRR
[…]
$\times$ 0.439 0.536 0.640 119 0.508
[…]
$\times$ 0.636 0.694 0.753 30 0.680
[…]
$\times$ 0.745 0.818 0.888 7 0.794
More Details for Experiments. For more details about the feasibility test, sensitivity experiment, loss functions, visualization results, and link evaluation, please refer to Appendix E.
5. Conclusion
In this paper, we introduce a new KGE model called KAAE. Specifically, to address the issue of semantic relevance, KAAE incorporates a joint interaction process that simultaneously captures the connections among entities, relations, and attributes within a unified framework. To improve semantic precision, we introduce a novel contrastive learning approach that enhances the diversity of training samples by optimizing ordinal samples based on cosine similarity and an entity-blending technique. Our KAAE model effectively enhances numerical reasoning in KGs. Extensive experiments on standard benchmarks such as US-Cities, Spotify, and Credit demonstrate KAAE’s effectiveness.
Appendix A Preliminaries
A.1. Knowledge Graph Embedding Methods
Our model is based on the existing KGE method to model and evaluate the plausibility of the triples. Considering the continuity and scheduling problem of the numeric values, in this paper, we use TransE(Bordes et al., 2013) and Order-embedding(Vendrov et al., 2016). $\bm{e_{h}}$,$\bm{e_{h}}$, and $\bm{e_{r}}$ denote the embedding of head entities, tail entities, and relations respectively.The reasonableness score of the triple $(\bm{e_{h},e_{r},e_{t}})$ is obtained as follows.
TransE. A translation-based method that was designed to optimize the formula $\bm{e_{h}+e_{r}\approx e_{t}}$
(17) $score(\bm{e_{h},e_{r},e_{t}})=\varepsilon-\|\bm{e_{h}+e_{r}-e_{t}}\|_{1/2},$
[…]
Order-embedding. Encoding the hierarchical and structural information contained in the numerical attributes of entities into the model with certain sequences to improve the model’s ability to process the data and get a better understanding of the semantics and structure of the knowledge. For numerical reasoning tasks, the top-ranked attributes usually have larger values(e.g.,80kg>70kg>60kg). To improve the efficiency and accuracy of the model, we propose the idea based on the score function of KGE as follows.
(18) $score(\bm{e_{h},e_{t}})=\varepsilon-\|max(0,\bm{e_{t}-e_{h}})\|^{2}$
where $\varepsilon$ is a hyperparameter.
Appendix B The Details of Proofs
B.1. The Loss of Contrastive Learning
We provide details of the derivations of KAAE in Eq. (12) and show that minimizing KAAE is approximately to minimize the combination of prediction and uniformity losses.
$\mathcal{L}_{CL}=$ $-\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}$
$\log\frac{\exp(\textbf{p}^{\top}\textbf{u}/\tau)}{\exp(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\mathcal{V}}\exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)}$
$=$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
$+\log(\exp(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\mathcal{V}}\exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau))$
(19) $\geq$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
$+\log\sum\nolimits_{v_{-}\in\mathcal{V}}\exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)$
$=$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
$+\log\sum\nolimits_{v_{-}\in\mathcal{V}}\exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)$
$=$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
$+\log\sum\nolimits_{v_{-}\in\mathcal{V}}\frac{\exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)}{|\mathcal{V}|}\cdot|\mathcal{V}|$
(20) $\geq$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\frac{\textbf{p}^{\top}\textbf{u}}{\tau}$
$+\log|\mathcal{V}|+\sum\nolimits_{v_{-}\in\mathcal{V}}\log\frac{\exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)}{|\mathcal{V}|}$
[…]
$+\frac{1}{|\mathcal{V}|}\sum\nolimits_{v_{-}\in\mathcal{V}}\log\exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)$
[…]
$+\frac{1}{|\mathcal{V}|}\sum\nolimits_{v_{-}\in\mathcal{V}}\textbf{v}^{\top}\textbf{v}_{-}$
[…]
$-\frac{1}{|\mathcal{V}|}\frac{1}{|\mathcal{V}|}\sum\nolimits_{v\in\mathcal{V}}\sum\nolimits_{v_{-}\in\mathcal{V}}\|\textbf{v}-\textbf{v}_{-}\|_{2}^{2},$
[…]
B.2. Proof of Theorem
Minimizing KAAE’s objective in Equation (12) with an exponential moving average is equivalent to maximizing mutual information between representation $V$ and the one-hop pattern $Y$:
(22) $\begin{split}\mathcal{L}_{CL}\geq H(V|Y)-H(V)=-I(V;Y),\end{split}$
Proof. According to the derivations in Appendix B.1, we have:
$\mathcal{L}_{CL}\geq$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\frac{1}{|\mathcal{N}(v)|}\sum\nolimits_{u\in\mathcal{N}(v)}-\textbf{p}^{\top}\textbf{u}$
$+\frac{1}{|\mathcal{V}|}\sum\nolimits_{v_{-}\in\mathcal{V}}\textbf{v}^{\top}\textbf{v}_{-}$
[…]
$-\frac{1}{|\mathcal{V}|}\sum\nolimits_{v\in\mathcal{V}}\sum\nolimits_{v_{-}\in\mathcal{V}}\|\textbf{v}-\textbf{v}_{-}\|_{2}^{2}$
(24) $=$ $\frac{1}{\mathcal{|V|}}\sum\nolimits_{v\in\mathcal{V}}\|\textbf{p}-\textbf{z}\|_{2}^{2}$
$-\frac{1}{|\mathcal{V}|}\frac{1}{|\mathcal{V}|}\sum\nolimits_{v\in\mathcal{V}}\sum\nolimits_{v_{-}\in\mathcal{V}}\|\textbf{v}-\textbf{v}_{-}\|_{2}^{2},$
where z $=\frac{1}{\mathcal{N}(v)}\sum\nolimits_{u\in\mathcal{N}(v)}$ u is the mean representations from one-hop neighborhoods of node $v$.Equation (23) holds because that gradient of u is being blocked in the target encoder with exponential moving average; thus it can be viewed as a constant that depends only on the target encoder but not the variable $f_{\theta}$. We can view the neighborhood representations $z$ as sampled from the conditional distribution given the pseudo label $Y$ which indicates the one-hop neighborhood pattern: $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$ because nodes belonging to the same semantic class tend to share similar one-hop neighborhood patterns. As demonstrated in [57], we can interpret the first term as a conditional cross-entropy between $V$ and another random variable $Z$ whose conditional distribution given the pseudo-label $Y$ indicates the one-hop neighborhood pattern: $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$:
[…]
(25) $=$ $H(V|Y)+\mathcal{D}_{KL}(V\|Z|Y)$
$\geq$ $H(V|Y),$
where $H(\cdot)$ and $\mathcal{D}_{KL}(\cdot)$ denote denote the entropy and KL-divergence, respectively. Minimizing the first term in Equation (24) is approximately minimizing $H(V|Y)$. We then inspect the second term in Equation (24). As shown in [58, 57], the second term is close to the differential entropy estimator:
(26) $\frac{1}{|\mathcal{V}|}\frac{1}{|\mathcal{V}|}\sum\nolimits_{v\in\mathcal{V}}\sum\nolimits_{v_{-}\in\mathcal{V}}\|\textbf{v}-\textbf{v}_{-}\|_{2}^{2}\varpropto H(V),$
As a result, minimizing the loss of KAAE can be seen as a proxy for maximizing the mutual information between the representations $V$ and the pseudo-labels $Y$ which indicates the neighborhood patterns of one hop nodes. Thus, the proof is completed.
Appendix C More Details of EXPERIMENTAL SETTINGS
The following subsections describe the data statistics, hyperparameters, experimental setup, and additional experimental results.
C.1. Dataset Statistics
To evaluate the performance of KAAE in this paper, we used the US-Cities, Spotify, and Credit datasets, and detailed information about the datasets is described in Table 7 below.
Baseline. We classified various existing representation learning methods into six groups: Euclidean KGE including TransE(Bordes et al., 2013), ConvE(Dettmers et al., 2018), ComplEx (Trouillon et al., 2016), TuckER(Balazevic et al., 2019b), DistMult(Dong et al., 2014) and HAKE(Zhang et al., 2020). Hyperbolic KGE including MuRP(Balazevic et al., 2019a), ConE(Bai et al., 2021), AttH(Chami et al., 2020) and GIE(Cao et al., 2022b); GNNs-KGE including R-GCN(Schlichtkrull et al., 2018) and WGCN (Shang et al., 2019); Attributed KGE including KBLRN(García-Durán and Niepert, 2018), MT-KGNN (Tay et al., 2017) and LiteralE(Kristiadi et al., 2019); and Self-supervised model for graph including BiGI(Cao et al., 2021), SLiCE(Wang et al., 2021a), SimGCL(Yu et al., 2022), and RAKGE(Kim et al., 2023); Deterministic stand-alone value representation methods including DICE(Sundararaman et al., 2020), NEKG(Duan et al., 2021) and NRN(Bai et al., 2023).
  • Euclidean KGE. A representation learning method that models entity and relation embeddings in Euclidean space. TransE(Bordes et al., 2013) uses Euclidean distance as a distance measure. ConvE(Dettmers et al., 2018), ComplEx (Trouillon et al., 2016), TuckER(Balazevic et al., 2019b), and DistMult(Dong et al., 2014) use inner product as a distance measure. HAKE (Zhang et al., 2020) capture hierarchical relations. At the same time, the performance improvement strategy of LTE-KGE(Zhang et al., 2022) was implemented in all Euclidean KGE baselines above.
  • Hyperbolic KGE. Mapping KGs to specific hyperbolic spaces. This method operates well in circumstances where relations in KGs are hierarchical. MuRP(Balazevic et al., 2019a), ConE(Bai et al., 2021), and AttH(Chami et al., 2020) model hierarchical relations in latent space. Furthermore, GIE(Cao et al., 2022b) considers both hierarchical and ring-like structures.
  • GNNs-KGE. A method that makes use of GCN layers as entity encoders and predicts the plausibility score of each triple using the score functions of other KGE models. In this paper, we select two well-performing baselines: R-GCN(Schlichtkrull et al., 2018) and WGCN (Shang et al., 2019).
  • Attributed KGE. Utilizing side information to obtain more robust embeddings in sparsely observed KGs. KBLRN(García-Durán and Niepert, 2018), MT-KGNN (Tay et al., 2017) and LiteralE(Kristiadi et al., 2019) are well-performing KGE models based on numerical attributes.
  • Self-supervised model for graph. An approach that enriches node representation through self-supervised learning. BiGI(Cao et al., 2021) and SLiCE(Wang et al., 2021a) capture both global and local structures to enhance node representation. SimGCL(Yu et al., 2022) adjusts the uniformity of the learned representation by adding random noise vectors to node representations. RAKGE(Kim et al., 2023) captures the connection between relations and numerical attributes to improve node representation.
  • Deterministic value representation methods. To validate the effectiveness of our numerical embedding method, we explore the use of deterministic stand-alone value representations. DICE(Sundararaman et al., 2020)creates corpus-independent deterministic embeddings for numbers. NEKG(Duan et al., 2021) embeds the knowledge graph in a vector space using a graph embedding method for obtaining embeddings of the number entities. NRN(Bai et al., 2023) introduce new numerical variables and operations to describe queries involving numerical attribute values, alternatively encoding entities and numerical values into separate encoding structures.
Table 7. Data statistics
Dataset US-Cities Spotify Credit
# of Triples All 60,231,493 6,098,734 4,712,388
Train 86,124 30,255 52,419
Valid 30,072,684 3,034,245 2,329,984
Test 30,072,685 3,034,245 2,329,985
# of Entities 9,296 3,374 1,790
# of Non-numeric Relations 3 3 3
# of Numeric Relations 3 3 3
# of Attributes 65 10 5
# of Numerical Triplets 465,505 20,170 8,915
C.2. Hyperparameters
To ensure the reproducibility of our model, we list the hyperparameters used in KAAE in Table8. These parameters were selected based on the validation splits across all three datasets: US-Cities, Spotify, and Credit.
Table 8. Hyperparameters to reproduce our results.
Hyperparameter US-Cities Spotify Credit
Entity/Relation Dimension $d_{emb}$ 200 200 200
Attribute Dimension $d_{att}$ 200 200 200
Number of Attention Heads L 5 5 5
Number of MoE Experts K 3 3 3
Score Function Bias $\varepsilon$ 9.0 9.0 9.0
Number of Top-K samples 3 3 3
Numerical Reasoning Weight $\Delta$ 0.25 0.25 0.25
CL Loss Coefficient $\lambda$ 0.25 0.25 0.25
Dropout Rate 0.7 0.7 0.7
Learning Rate 0.001 0.001 0.001
Batch Size 200 256 256
Appendix D LEARNING PROCEDURE
The overall learning process of KAAE is shown in Algorithm 1. At the beginning, we initialize the loss. We then use the KA Encoder to obtain the header embedding $\bm{e_{joint,h}^{att}}$ for each triple(h, r, t) in the training set (described in line 5). If r is a numerical comparison relation, we obtain the set $\mathcal{E}^{+}$, $\mathcal{E}^{-}$ of comparative positive and negative samples based on $\bm{e_{joint,h}^{att}}$ (described in lines 7-8). Lines 9-10 show the process of generating high-quality blended positive/negative samples via Top-K sampling. Then, the contrast learning loss is accounted for in the total loss (line 11). Finally, the cross-entropy loss is added according to Eq(14), and finish the updating process.
Algorithm 1 Training KAAE
1:  Input: Training set $\mathcal{T}=\mathcal{G}\cup\mathcal{G}^{-}$; Positive sample set for $\mathcal{P}[e,r]=\{i|(e,r,i)\in\mathcal{G}\}$; Negative sample set for $\mathcal{N}[e,r]=\{j|(j,r,e)\in\mathcal{G}\}$; Initialized MoEKA_Encoder; Comparison relation set $\mathcal{R}_{Comparison}$; Coefficient for contrastive loss $\lambda$; # of epochs for training epochs
2:  for  1,2$\cdots$,epochs do
3:     Initialize loss L = 0
4:     for each (h, r, t) $\in\mathcal{T}$ do
5:        Acquire the head embedding $\bm{e}_{joint,h}^{att}$ = MoEKA_Encoder(h, r)
6:        if r $\in\mathcal{R}_{Comparison}$ and (h, r, t) $\in\mathcal{G}$ then
7:           Get $\mathcal{E}^{+}$ = {MoEKA_Encoder(p, r)p $\in\mathcal{P}[h,r]$}
8:           Get $\mathcal{E}^{-}$ = {MoEKA_Encoder(n, r)n $\in\mathcal{N}[h,r]$}
9:           Generate a high-quality blending positive $\bm{e}_{mix}^{t+}$ with $\mathcal{E}^{+}$ and $\bm{e}_{joint,h}^{att}$ via Eqs. (9),(10)
10:           Generate a high-quality blending negative $\bm{e}_{mix}^{t-}$ with $\mathcal{E}^{-}$ and $\bm{e}_{joint,h}^{att}$ via Eqs. (9),(11)
11:           $\mathcal{L}=\mathcal{L}$ + $\lambda\mathcal{L}_{CL}$,($\mathcal{L}_{CL}$ via Eq.(12))
12:        end if
13:        $\mathcal{L}=\mathcal{L}$ + $\mathcal{L}_{BCE}$,($\mathcal{L}_{BCE}$ via Eq.(14))
14:     end for
15:     Update
16:  end for
Appendix E ADDITIONAL EXPERIMENTAL RESULTS
E.1. Feasibility Test
To test the ability of KAAE to learn values based on the magnitude of attribute numerical values, we conducted a feasibility experiment. We convert the values in the Spotify dataset to binary values to distinguish the presence or absence of numeric attributes. We assign a value of 1 to data that exists and a value of 0 to data that does not exist. The result is shown in Table 9. Our proposed model KAAE outperforms the competitors LiteralE(Kristiadi et al., 2019) and RAKGE(Kim et al., 2023). LiteralE did not show significant changes under the two experimental conditions. Both KAAE and RAKGE can compare the magnitude of the values, and KAAE performed better.
Table 9. Feasiblity test on Spotify dataset
Model Setting H@1 H@3 H@10 MR MRR
LiteralE Binary 0.263 0.362 0.480 79 0.339
Numeric 0.266 0.377 0.494 74 0.347
RAKGE Binary 0.394 0.479 0.571 62 0.459
Numeric 0.487 0.602 0.692 37 0.560
KAAE Binary 0.474 0.588 0.683 47 0.550
Numeric 0.636 0.694 0.758 30 0.680
E.2. Sensitivity Experiment
We adjust the hyperparameter $\lambda$ of contrastive learning loss in Eq. (15) to verify the necessity of introducing contrastive learning. At the same time, we changed $\Delta$ in Eq. (13) to test the performance of KAAE with different weights of order-embedding in the score function. The results are shown in Figure 6.
Figure 6. Performance of KAAE with different hyperparameter values of $\lambda$ and $\Delta$ on Credit dataset.
E.3. Loss Experiment
To verify the effectiveness of our proposed two loss functions, we compare the performance of KAAE trained with different loss functions. The results in Table 10 show that both loss functions are reasonable.
(27) $\begin{split}{Loss}_{1}=-\frac{1}{\mathcal{|G|}}\sum\limits_{l\in\mathcal{|G|}}\log\sigma(&score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{mix}^{t+})\\ -&score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{mix}^{t-})),\end{split}$
(28) $\begin{split}{Loss}_{2}=-&\frac{1}{\mathcal{|G|}}\sum\nolimits_{l\in\mathcal{G}}\log\frac{exp(\textbf{p}^{\top}\textbf{u}/\tau)}{exp(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\bm{e}_{mix}^{t-}}exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)},\end{split}$
Table 10. Loss function test
Dataset Loss H@1 H@3 H@10 MR MRR
US-Cities Loss1 0.441 0.532 0.645 121 0.511
Loss2 0.439 0.536 0.640 119 0.508
Spotify Loss1 0.640 0.696 0.758 34 0.682
Loss2 0.636 0.694 0.753 30 0.680
Credit Loss1 0.732 0.811 0.890 8 0.786
Loss2 0.745 0.818 0.888 7 0.794
E.4. Visualization
Importance of entity-aware The results of the experiment to observe the effect of entity information on relation-attribute attention scores are shown in Figure 7, indicating that adding information from an entity to each relation effectively improves the problem of equalizing the attention to attributes in some relations. For example, when predicting the relation LIMIT_BILL_comp, the proximity of the MoEKA Encoder’s attention to the five attributes results in an inability to capture the focus, and additional consideration of information from entity ID1763 results in the MoEKA Encoder placing more emphasis on the LIMIT_BALL and TOTAL_BILL attributes.
Figure 7. Visualization of relevance scores of each numeric attribute without/with entity-aware on Credit dataset.
Analysis of Contrastive Learning. To improve the continuity of numerical attributes and the sensitivity of numerical values to very small differences near boundary values, we propose a method to accomplish contrastive learning using ordinal high-quality positive and negative samples. We use the algorithm named T-SNE(Maaten and Hinton, 2008) to visualize the samples in 2D space to observe the distribution of high-quality positive and negative samples. First, we replace the vector $\bm{e_{joint,i}^{att}}$ in Eq.(7), which contains rich information of numerical attribute and order-embedding. Eqs. (10)-(11) are repeated 100 times using a definite attribute value of age-40 with entity and relation to accumulating a large number of tail samples in contrastive learning. We compare KAAE to another well-performed model, RAKGE(Kim et al., 2023). The results in Fig.8 show that our KAAE can improve the generalization ability and robustness of the model by learning difficult samples to get a good representation of entities and make more accurate judgments in numerical reasoning tasks.
(a) RAKGE on US-Cities dataset
(b) RAKGE on Spotify dataset
(c) RAKGE on Credit dataset
(d) KAAE on US-Cities dataset
(e) KAAE on Spotify dataset
(f) KAAE on Credit dataset
Figure 8. Visualization of the entity embeddings using T-SNE and K-means. The first row shows the results of RAKGE on the three datasets and the second row shows the better performance of KAAE. By using KAAE, the nodes are more compact with each other, which means that we can get excellent entity embeddings
E.5. Evaluation on FB15k-237 Dataset
FB15k-237 is a classical dataset used for the standard link prediction task accomplished by knowledge graphs. Therefore, we conducted additional experiments on the FB15k-237 dataset. As we can see from Table11, KAAE is still the state-of-the-art model.
Table 11. Link prediction results on FB15k-237(Toutanova and Chen, 2015). Bold scores represent the best results.
Model H@1 H@3 H@10 MR MRR
TransE(Bordes et al., 2013) 0.240 0.368 0.516 182 0.332
DisMult(Dong et al., 2014) 0.202 0.306 0.433 392 0.279
ConvE(Dettmers et al., 2018) 0.232 0.351 0.492 276 0.319
TuckER(Balazevic et al., 2019b) 0.266 0.394 0.544 - 0.358
MuRP(Balazevic et al., 2019a) 0.245 0.370 0.521 - 0.336
R-GCN(Schlichtkrull et al., 2018) 0.153 0.258 0.414 - 0.248
LiterlE(Kristiadi et al., 2019) 0.232 0.348 0.483 280 0.317
KBLRN(García-Durán and Niepert, 2018) 0.215 0.333 0.468 358 0.301
MT-KGNN(Tay et al., 2017) 0.204 0.312 0.445 532 0.285
RAKGE(Kim et al., 2023) 0.282 0.402 0.531 321 0.366
KAAE 0.355 0.464 0.553 164 0.428
E.6. Analysis of Backbones
To validate the compatibility of TransE with KAAE, we replaced different backbones. Specificaly, we changed TransE to RotatE(Sun et al., 2019), SimplE(Kazemi and Poole, 2018) and DistMult. Table12 (in Appendix) shows that our KAAE+TransE model significantly outperforms the KAAE+RotatE model, exhibiting enhancements of 354.2% in Hits@1 and 590.4% in MRR. It also outperforms SimplE and DistMult. We infer that this is because TransE(Bordes et al., 2013) is compatible with sequential embedding, which helps the model to solve the ordering and boundary problems of values and deal with asymmetric relationships. In this sense, replacing the backbones causes the model to have difficulty retaining the ordering information.
E.7. Breakdown Analysis.
Based on our current framework, we have selected some of the relations that are difficult to recognize. In dataset US-Cities, the difference in the values of each attribute for each city is almost always less than 1, which makes it difficult to determine the relation is_located_in_county and is_located_in_state. This gives us ideas for improvements. In our future work, we will explore how to target the improvement of the link prediction tasks involving fine-grained numerical size comparison relations. In addition, we are already working the challenges of extreme values on model robustness.
Table 12. Backbone study of KAAE on Credit dataset
Model H@1 H@3 H@10 MR MRR
KAAE+TransE 0.745 0.818 0.888 7 0.794
KAAE+RotatE 0.164 0.116 0.205 151 0.115
KAAE+SimplE 0.026 0.201 0.447 39 0.163
KAAE+DistMult 0.003 0.080 0.200 177 0.070
Relevance of Attributes. Figure 9 illustrates the change in performance of KAAE for different densities of numerical attributes. As the data density decreases, LiteralE, a model that relies on structural information for numerical reasoning, doesn’t change much, at the same time the performance of our KAAE model decreases, but always stays ahead in absolute performance. We can infer that through our MoEKA Encoder and learning by utilizing high-quality positive and negative samples in contrastive learning, KAAE is sensitive to numerical magnitude while possessing the ability to learn adaptively and remain robust despite sparse numerical attributes.
Figure 9. Performance of KAAE with different density of numeric attributes for Spotify dataset. A density of 100% means that all entities have numeric values for every attribute, while 0% means that none of the entities have numeric attributes.